Q) Price of petrol was Rs. 60 per litre in 2015 and Rs. 100 per litre in 2021. The average wage in manufacturing was Rs. 240 per hour in 2015 and 300 in 2021.

a.     By what percentage did the price of a litre of petrol rise? [3 marks]

b.    By what percentage did the wage rise? [3 marks]

c.     In each year, how many hours does a worker have to work to earn enough to buy 10 litres of petrol? [3 marks]

d.     Did workers’ purchasing power in terms of petrol rise or fall? [3 marks]



1
Expert's answer
2021-06-16T09:22:49-0400

a) percentage rise in price of petrol in 2021 in comparison to 2015"=(100-\frac{60}{60})\times100= 66.67\%"

b) Percentage rise in wages in 2021in comparison to 2015"=( 300-\frac{240}{240})\times100= 25\%."

c)In 2015:amount of work for 10l petrol="10\times 60=Rs .600"

working hours will be

working hours=amount for 10l petrol "\div" average wage in 2015

"=\frac{600}{240}=2.5 hours"

for year 2021

amount of work for 10l petrol="10\times 100=Rs .1000"

working hours will be

working hours=amount for 10l petrol "\div" average wage in 2021

"=\frac{1000}{300}=3.3 hours"

d:workers’ purchasing power in terms of petrol fall